### v4

@@ Abstract
In this paper, I develop and characterize a random attention model with unobserved reference alternatives. The decision-maker pays attention to different subsets of the available set of alternatives randomly. The reference alternatives are exactly those alternatives that are always paid attention to, i.e. they are attention-privileged. These alternatives are unknown to the outside observer. The characterization allows for a complete identification of the reference alternatives and a coarse identification of the underlying preferences. I then restrict the model by considering the independent random attention function and provide a complete identification of the underlying preferences.
@@ 1 Introduction
Decision-making is often constrained by limited attention, whether due to cognitive processing limits caused by the availability of too many alternatives or a shortage of time. In such situations, decision-makers (DM) often resort to heuristics that can help them decide relatively quickly with less mental exercise. If the DM has to make similar choices multiple times with the same underlying primitives, the heuristic that decides how to allocate attention can lead to different choices. Stochastic choice functions represent such choices, with the main research task then being to ascertain the likely underlying heuristic and the primitives that drive the decisions of the DM. Cattaneo et al. (2020) introduced the Random Attention Model (RAM) framework to deal with the analysis of stochastic choice functions with random choice due to limited attention.
@@
A substantial body of literature has explored how an observable reference affects the choices of a DM. In particular, Kovach and Suleymanov (2023) uses the framework of RAM to look at how a reference alternative, or a status quo in their case, can affect choices if it directs attention. They were the first to deal with observable reference alternatives’ impact on attention in a stochastic choice setting. It is reasonable to assume that a reference alternative can direct attention when the underlying heuristic is applied due to limited attention. They provide a characterization of a reference-dependent RAM, as well as special cases where the attention function is further restricted.
@@
In certain cases, it is logical to assume that the reference is observable. For instance, for a DM in the market for a new car, her current car can be a reference. However, in many situations, it is not realistic to assume the knowledge of the DM’s reference to anyone else but the DM themselves. Further, to analyze a choice problem with observable references, we have to deal with a family of stochastic choice functions, as done by Kovach and Suleymanov (2023). The inputs for the stochastic choice functions in their model were pairs of subsets [^3] Hereafter referred to as menus. and references- one family of stochastic choice function for each alternative as the reference. This is quite a significant data requirement. If we restrict the references to be unknown, we deal with just a single stochastic choice function over menus and thus, we do not have to deal with families of stochastic choice functions for each possible reference.
It is also not obvious that a reference alternative must be unique. In what follows, I allow for multiple reference alternatives; clearly, one particular such alternative could be the status quo. In this spirit, I define an alternative to be a reference alternative if and only if it is attention privileged, i.e. it is always paid attention to. Since I am looking at choices from menus that depend on unobservable reference alternatives, I assume that the references are menu-independent. It is certainly plausible that references can be menu-dependent, through which we can explain a larger set of stochastic choice behavior, but dealing with menu-dependent unobservable references is beyond the scope of this paper. I provide below four examples that match the mentioned assumptions.
@@
  • Example 1: University Choice: High School students making a choice of university to continue their higher education in the UK filter out the large number of choices into smaller consideration sets. The universities are divided into old and new universities. This exogenous distinction impacts what universities are considered by students. For instance, Dawes and Brown (2005) found that minority ethnic groups had a greater proportion of old universities in their consideration set, possibly due to reputation concerns of these ethnic groups. The college choice of high school students in the US is also a pretty involved process. Rosen et al. (2008) showed that what major is offered as well as location were crucial factors in shortlisting colleges. Parental and guidance counselor advice also has a significant impact on the decision-making process. It is conceivable that colleges that are close by act as references. Dawes and Brown (2005) observed a similar phenomenon in the UK data as well. If the guidance of a counselor or a parent impacts which colleges act as references, it is not possible to observe these references.
@@
  • Example 2: Apparel Shopping: While shopping for apparel, brand name is often crucial in drawing attention to particular items. The Air Jordan always catches the eye due to its association with Michael Jordan. However, though brand awareness might be enough to pull attention, it is not enough to provide a positive brand association as found by Dew and Kwon (2009). Thus, attention and preference are independent attributes.
@@
  • Example 3: Political Choice: A stochastic choice framework can also be used to model a population of voters with similar preferences but with access to different options. A nationally strong party (a reference) will always be considered; however, a weaker party in one district might be a stronger party in another and be considered there, leading to different choices even when voters exhibit homogenous preferences. Wilson (2008) showed this for the Mexico national elections in 2000. The biggest party in Mexico, the Partido Revolucionario Institucional (PRI) has a strong presence in all the states, which makes it the reference, while the two opposition parties have varied presence in different states.
@@
  • Example $4$I would like to thank Marco Mariotti for this interpretation of the model.: Suppose past choices of the DM influence their future choices, i.e. they act as a reference. The outside observer only observes the current feasible sets and the choices from those but not the choices in the past. In this case, (multiple) past choices act as references and are thus, unobservable to the researcher. If the reference depends on recommendations from advisors (college choice) or past choices in elections (political choice), it is possible that the researcher is not able to observe the reference.
@@
In this paper, I look at a random attention model with unobservable menu-independent reference alternatives that are attention-privileged (RAM-UR). I provide a characterization of the RAM-UR by using three axioms and identifying the corresponding primitives (i.e., the set of reference alternatives and the underlying preferences). One can uniquely identify the set of references but the underlying preferences are only coarsely identified. The identification of the preferences gives the relative ranking between any alternative and any reference alternative and the transitive closure thereof. I then restrict the attention function to be of the independent random attention (IRA) form (Manzini and Mariotti (2014)) and show that under such a restriction, the preferences can be fully identified. The existence of menus with no references is crucial in making the preference identification in this case. This is possible because of the menu-independence of the references. Kovach and Suleymanov (2023) allowed for the references to be part of the primitive which precludes them from considering menus without references. I also elaborate on the fact that rational choice is a boundary case of both the RAM-UR and the RAM-UR-IRA, while the IRA model is itself a boundary case of the RAM-UR-IRA [^5] Boundary cases can be of two types: When the set of references is the set of all alternatives, or when the set of references is empty..
@@
The rest of the paper is organized as follows. In Section 2, I introduce the general RAM-UR model, the axiomatic characterization of the model is presented in Section 3 along with a discussion on the identification results. The introduction and characterization of the RAM-UR-IRA are then provided in Section 4, where I also discuss the nestedness properties of the corresponding random attention models. Section 5 relates my work to the existing literature. All proofs are relegated to Appendix 1.
2 Random Attention Model with Unobserved References
@@
Let $X$ be the finite set of all possible alternatives. Denote the set of all subsets of $X$ by $\mathcal{X}$, with a typical element being $A$.
@@
Below I denote by $a^{*}$ the default alternative and interpret it as abstention or failing to make a choice or delaying choice. The idea of default is common in the literature on attention, especially the literature originating from Manzini and Mariotti (2014). The assumption is that DMs prefer to get something over abstention [^6] Equivalently, the default is only chosen if none of the alternatives is paid attention to.. For each $A\in\mathcal{X}\setminus\phi$, define $p(a^{*},A)=1-\sum_{a\in A}p(a,A)$ to be probability of abstention from the menu $A$. I will also denote for any $A\in\mathcal{X}\setminus\phi$, $A^{*}=A\cup\{a^{*}\}$.
@@
Here $\mu(D,A)$ can be interpreted as the probability that $D$ is the consideration set when the menu offered is $A$. It is natural that if $D$ is not a subset of $A$, then the function should take a value of zero. The second point also stresses the fact that each subset of $A$ can be a consideration set with some positive probability (full support assumption). In contrast to the above definition, Cattaneo et al. (2020) impose a monotonicity condition on the attention function $\mu$ and do not impose the full support assumption. Kovach and Suleymanov (2023) imposed the full support assumption and it will be crucial in our model as well to identify the set of references. Without the full support assumption, we cannot differentiate between the reference and some of the non-reference alternatives. I also relax the monotonicity condition in the general model. As demonstrated later, this relaxation does not result in much loss in identification. A random attention model (RAM) is then defined as follows.
@@
Since the empty set is also a possible consideration set for each menu, the probability of choosing the default is just the probability of the empty set being the consideration set. Note that the RAM without the full support assumption has no empirical content since we can just attribute the probability of each alternative in a menu to the singleton menu containing that alternative and the probability of abstention to the attention probability of the empty set. The empirical content of the RAM in Cattaneo et al. (2020) comes from the monotonicity restriction they impose on the attention function.
Below I introduce the Random Attention Model with Unobserved References with references being exactly those alternatives that are attention privileged.

### v5

@@ Abstract
In this paper, I develop and characterize two models of random attention that differ from each other with respect to the menu-dependence of the unobserved reference alternatives. In both models, the decision-maker pays attention to subsets of the available set of alternatives randomly with the reference alternatives being always paid attention to. Under menu-dependence, partial identification of both the reference alternatives and the underlying preferences is provided. For the case of multiple menu-independent references, I provide a complete identification of the references and a coarse identification of the underlying preferences. A complete identification of the latter is provided when the independent random attention function is considered.
@@ 1 Introduction
Decision-making is often constrained by limited attention, whether due to cognitive processing limits caused by the availability of too many alternatives or to a shortage of time. In such situations, decision-makers (DM) often resort to heuristics that can help them decide relatively quickly with less mental exercise. If the DM has to make similar choices multiple times with the same underlying primitives, the heuristic that decides how to allocate attention can lead to different choices. Stochastic choice functions represent such choices, with the main research task then being to ascertain the likely underlying heuristic and the primitives that drive the decisions of the DM. Cattaneo et al. (2020) introduced the Random Attention Model (RAM) framework to deal with the analysis of stochastic choice functions with random choice due to limited attention.
@@
A substantial body of literature has explored how an observable reference affects the choices of a DM. In particular, Kovach and Suleymanov (2023) uses the framework of RAM to look at how a reference alternative, or a status quo in their case, can affect choices if it directs attention. These authors were the first to deal with observable reference alternatives’ impact on attention in a stochastic choice setting. They provide a characterization of a reference-dependent RAM, as well as special cases where the attention function is further restricted.
@@
In certain cases, it is logical to assume that the reference is observable. For instance, for a DM in the market for a new car, her current car can be a reference. However, in many situations, it is not realistic to assume the knowledge of the DM’s reference to anyone else but the DM themselves. Further, to analyze a choice problem with observable references, we have to deal with a family of stochastic choice functions, as done by Kovach and Suleymanov (2023). The inputs for the stochastic choice functions in their model were pairs of subsets [^3] Hereafter referred to as menus. and references- one family of stochastic choice function for each alternative as the reference. This is quite a significant data requirement. If we restrict the references to be unknown, we deal with just a single stochastic choice function over menus.
It is also not obvious that a reference alternative must be unique. In one of the models presented in this paper, I allow for multiple reference alternatives; clearly, one particular such alternative could be the status quo. In this spirit, I define an alternative to be a reference alternative if and only if it is attention privileged, i.e. it is always paid attention to. The first model I look at has for each menu, an unobservable reference alternative that directs attention, and each menu can potentially have any alternative as the reference: the general model makes no restriction on how the reference changes with changing menus.
The second model allows for multiple reference alternatives. In particular, it defines multiple menu-independent references, i.e., there is a set of alternatives that are always attention privileged whenever offered. I provide four examples below that relate to the assumptions that I make in this paper.
@@
  • Example 1: University Choice: High School students making a choice of university to continue their higher education in the UK filter out the large number of choices into smaller consideration sets. The universities are divided into old and new. This exogenous distinction impacts what universities are considered by students. For instance, Dawes and Brown (2005) found that minority ethnic groups had a greater proportion of old universities in their consideration set, possibly due to reputation concerns of these ethnic groups. The college choice of high school students in the US is also a rather elaborate process. Rosen et al. (2008) showed that what major is offered as well as location, are crucial factors in shortlisting colleges. Parental and guidance counselor advice also has a significant impact on the decision-making process. It is conceivable that colleges that are close by act as references. Given the choice of colleges, the reference then changes depending upon the menu. Dawes and Brown (2005) observed a similar phenomenon in the UK data as well. If the guidance of a counselor or a parent impacts which colleges act as references, it is not possible to observe these references.
@@
  • Example 2: Apparel Shopping: Brand name is often crucial in drawing attention to particular items when it comes to shopping for apparel. The Air Jordan always catches the eye due to its association with Michael Jordan. However, though brand awareness might be enough to pull attention, it is not enough to provide a positive brand association as found by Dew and Kwon (2009). Thus, attention and preference seem to be independent attributes. If attention is brand-dependent, it would make sense that attention privilege is menu-independent: Air Jordans are always paid attention to when available.
@@
  • Example 3: Political Choice: A stochastic choice framework can also be used to model a population of voters with similar preferences but with access to different options. A nationally strong party will always be considered as a reference; however, a weaker party in one district might be a stronger party in another and be considered there, leading to different choices even when voters exhibit homogenous preferences. So, some alternatives might be references whenever available, while some alternatives’ being references are context dependent. This observation has been supported by Wilson (2008)’s analysis of the Mexico national elections in 2000. The biggest party in Mexico, the Partido Revolucionario Institucional (PRI) has a strong presence in all the states, which makes it the reference, while the two opposition parties have varied presence in different states.
@@
  • Example 4: Suppose past choices of the DM influence their future choices, i.e. they act as a reference. An outside observer can have information about the current feasible sets and the choices from those, but not the choices in the past. In this case, (multiple) past choices act as references and thus, they are unobservable for everyone but the DM. If the reference depends on recommendations from advisors (college choice) or past choices in elections (political choice), it is possible that the researcher is not able to observe the reference. [^4] I thank Marco Mariotti for providing such an interpretation of the model.
@@
In this paper, I look at two generalizations of the random attention model: one where each menu is endowed with an unobservable reference which is attention privileged (GRAM-UR); and another with unobservable menu-independent, possibly multiple reference alternatives that are attention-privileged (RAM-UR). I provide a characterization of the GRAM-UR and the RAM-UR by using testable axioms and identify the corresponding primitives (i.e., the set of reference alternatives and the underlying preferences). In the GRAM-UR, multiple representations are possible and references change depending upon the representation; however, one can uniquely identify the set of references for the RAM-UR, but the underlying preferences are only coarsely identified. I then restrict the attention function to be of the independent random attention (IRA) form (Manzini and Mariotti (2014)) and show that under such a restriction, the preferences can be uniquely identified. The existence of menus with no references is crucial in making the preference identification in this case. This is possible because of the menu independence of the references. Kovach and Suleymanov (2023) allowed for the references to be part of the primitive, which precludes them from considering menus without references [^5] In the GRAM-UR, menus without references are not possible, but they are possible in the RAM-UR and the RAM-UR-IRA.. It is also notable that rational choice is a boundary case of the RAM-UR and the RAM-UR-IRA, while the IRA model is itself a boundary case of the RAM-UR-IRA [^6] Boundary cases can be of two types the RAM-UR or RAM-UR-IRA: When the set of references is the set of all alternatives, or when the set of references is empty.. Rational choice is also a special case of the GRAM-UR.
@@
The rest of the paper is organized as follows. In Section 2, I introduce the generalized RAM-UR (GRAM-UR) model, the axiomatic characterization along with a discussion on the identification results. Section 3 introduces and characterizes the RAM-UR and discusses the identification results, compares them to the GRAM-UR and past models of limited attention. The introduction and characterization of the RAM-UR-IRA are then provided in Section 4, where I also discuss the nestedness properties of the corresponding random attention models. Section 5 relates my work to the existing literature. All proofs are relegated to Appendix 1.
2 Generalized Random Attention Model with Unobserved References
@@
Let $X$ be the finite set of all possible alternatives. Denote the set of all subsets of $X$ by $\mathcal{X}$, with a typical element denoted by $A$.
@@
The third part of the above definition allows for the sum of choice probabilities of alternatives from a menu not to sum up to one. I denote by $a^{*}$ the default alternative and interpret it as abstention, failing to make a choice or delaying choice. The idea of default is common in the literature on attention, especially the literature originating from Manzini and Mariotti (2014). The assumption is that DMs prefer to get something over abstention [^7] Equivalently, the default is only chosen if none of the alternatives is paid attention to.. For each $A\in\mathcal{X}\setminus\phi$, define $p(a^{*},A)=1-\sum_{a\in A}p(a,A)$ to be probability of abstention from the menu $A$. I will also denote for any $\in\mathcal{X}\setminus\phi, A=A$$^{*}=A\cup\{a^{*}\}$. The inclusion of the default in a choice model is not ideal, however, since the probability of default is just the residual probability in case the observable choice probabilities do not sum up to one, analysis is possible if we restrict the model to have probabilities sum up to one; we do that in GRAM-UR.
@@
Here $\mu(D,A)$ can be interpreted as the probability that $D$ is the consideration set when the menu offered is $A$. It is natural that if $D$ is not a subset of $A$, then the function should take a value of zero. The second point also stresses the fact that each subset of $A$ can be a consideration set with some positive probability (full support assumption). In contrast to the above definition, Cattaneo et al. (2020) impose a monotonicity condition on the attention function and do not impose the full support assumption. Kovach and Suleymanov (2023) imposed the full support assumption. It will be crucial in our model as well to identify the set of references. A modified full support assumption in line with the idea that references are exactly those alternatives that are attention-privileged is quite natural. I also relax the monotonicity condition in my models. A random attention model (RAM) is then defined as follows.
@@
Since the empty set is also a possible consideration set for each menu, the probability of choosing the default is just the probability of the empty set being the consideration set. Note that the RAM as defined above without the full support assumption has no empirical content since we can just attribute the probability of each alternative in a menu to the singleton menu containing that alternative and the probability of abstention to the attention probability of the empty set.
In the generalized random attention model with unobserved references, each menu comes endowed with a reference alternative, which is attention-privileged. The attention function then takes the following form:
Definition 4
Let $r_{A}\in A\subseteq X$ denote the reference alternative for the menu $A$. A generalized reference-dependent attention function is a mapping $\mu_{r}:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$ such that:
  1. 1.
    $\sum_{D\subseteq A}\mu_{r}(D,A)=1$.
  2. 2.
    $\mu_{r}(D,A)>0$ if and only if $\{r_{A}\}\subseteq D\subseteq A$ [^9] This is the modified full support assumption, every subset which maintains the attention privilege of the reference is paid attention to with positive probability..
The idea here is similar to attention privilege as described by Kovach and Suleymanov (2023), i.e. a reference is always considered. In fact, that is how a reference is defined in this setting. An alternative is a reference if and only if it is attention privileged. This definition can be extended to allow for multiple references, which we do in the model in section 3. Note that if there are no references in any menu, this definition is the same as a standard attention function as defined in Definition 2.
The modified full support assumption says that inherently, there is no bias towards any set when it comes to attention, apart from the bias to always consider the reference. The reference being menu-dependent means that the reference itself could be context-dependent: for instance, the biggest item could be the reference amongst any set of alternatives, which means that the reference keeps changing based on the set. However, I do not impose any consistency requirement on the reference across menus in the general model.
The fundamental analytical task with respect to the presented model is to identify the reference alternative and the underlying preferences. In most canonical papers, it is assumed that the references are observable, which precludes the need to identify the set of references. In the RAM model introduced by Cattaneo et al. (2020), the monotonicity condition is crucial in making a partial identification of the underlying preferences. The (partial) identification of the preferences in the current model is aided by the full support assumption and the presence of attention-privileged alternatives.
Definition 5
A stochastic choice function $p$ is a generalized random attention model with unobserved references (GRAM-UR) if for all menus there exists a reference alternative $(r_{A})_{A\subseteq X}$, a strict total order $\succ$ over $X$ and a generalized reference-dependent attention function $\mu_{r}:\mathcal{X}\times\mathcal{X}\rightarrow[0,1]$ such that for all $A\in\mathcal{X}\setminus\phi$ and $x\in A$:
$p(x,A)=\sum_{D\in\{D\in\mathcal{X}:x\succ y,\forall y\in D\setminus\{x\}\}}\mu_{r}(D,A)$
The above definition is similar to the definition of RAM, apart from the fact that it allows for a generalized reference-dependent attention function.
A feature of the GRAM-UR is that the default alternative is never chosen with a positive probability in any menu. This is because the reference acts as the de facto default, i.e. whenever one of the other alternatives is not chosen, the DM sticks with the default. This property of the reference is similar to its function as a status quo, as modeled by several papers like Masatlioglu and Ok (2005) and Masatlioglu and Ok (2014).
The partial identification of underlying preferences relies on the fact that any alternative that is not chosen with positive probability in a menu is not preferred to the attention-privileged reference, and any alternative that is chosen with a positive probability beats the reference. In the next section, I describe an axiomatic characterization of the GRAM-UR using testable axioms that ensure consistency of the preferences elicited in the way described above [^10] The elicited preferences are inconsistent if they violate asymmetry or transitivity..
2.1 Characterization
In this section, I characterize the GRAM-UR using two testable axioms.
Acyclicity of No Choice (ANC): Consider a sequence of alternatives $x_{1},x_{2},...,x_{k}\in X$, $k\geq 2$ such that there exist $A_{1},A_{2},...,A_{k-1}\in\mathcal{X}$ with $p(x_{i},A_{i})>0$ and $p(x_{i+1},A_{i})=0$ for all $i\in\{1,2,...,k-1\}$, then there does not exist $A_{k}\in\mathcal{X}$ such that $p(x_{k},A_{k})>0$ and $p(x_{1},A_{k})=0$.
This is an axiom similar to one imposed by Cattaneo et al. (2020) to characterize the RAM. It disciplines the way a probability-zero choice is observed. In the GRAM-UR, the full support condition implies that zero probability choice for an alternative signifies that the reference is preferred to the alternative. In the same menu, if another alternative is chosen with a positive probability, then the reference cannot be preferred over it. ANC ensures we do not observe a cycle of such probabilities.
No Default (ND): For each $A\in\mathcal{X}$, $\sum_{x\in A}p(x,A)=1$.
The ND condition is usually assumed in the primitives of choice models. I have introduced it as an axiom instead to allow the GRAM-UR to lie in the same framework as other attention models like that of Manzini and Mariotti (2014) and the other models I introduce later in the paper.
Theorem 1
A stochastic choice function is representable by a generalized random attention model with unobserved references if and only if it satisfies ANC and ND.
Note that if a stochastic choice function has a GRAM-UR representation, such a representation is not unique. The proof in Appendix 1 shows that there are possible extensions of the partially identified preferences (Szpilrajn (1930)) each of which provides a representation. However, when an extension is fixed, the reference for each menu is uniquely identified. Given that we are working with unobservable references and significant restrictions to the primitives as compared to Kovach and Suleymanov (2023), it is natural that the preference identification result is weaker. However, it is not weaker than the identification results of the RAM by Cattaneo et al. (2020).
2.2 Multiple References Interpretation
A possible interpretation of the GRAM-UR model is for each menu to have multiple reference alternatives that are attention-privileged. However, in such a case, the most preferred reference alternative in any menu is the de facto default, which prevents any of the other references to be chosen with positive probability. This means that one possible representation for a GRAM-UR could be where all alternatives that are chosen with zero probability and the lowest preferred alternative (according to the extension) are all references. This representation of the GRAM-UR incorporates the idea of multiple reference alternatives. The attention function associated with the multiple reference interpretation takes interesting forms. For instance, the function that maps each menu to the set of references it contains could satisfy the attention filter by Masatlioglu et al. (2012), or the competition filter by Lleras et al. (2017). For the example below, partial identification of preferences as well as the references is possible.
Example 5
Consider the set of alternatives $X=\{a,b,c\}$ and the following observed choice probabilities:
$p(c,\{a,b,c\})=p(b,\{a,b,c\})=1/2$
$p(b,\{b,c\})=1\hskip 14.22636pt$
Here, the choice from $\{b,c\}$ gives us that $b$ is the reference in that menu and that $b$ is preferred to $c$. Since $a$ is not chosen with a positive probability in $\{a,b,c\}$, we have that $c$ is preferred to $a$. Here, in fact, we can completely and uniquely identify the underlying preferences: $b\succ c\succ a$ with the reference in each set being: $\{a,b\}\rightarrow a,\{a,c\}\rightarrow a,\{b,c\}\rightarrow b,\{a,b,c\}\rightarrow c$.
The multiple references interpretation allows that in $\{b,c\}$ the set of references is $\{b,c\}$ and in $\{a,b,c\}$ the set of references is $\{a,c\}$.
In the next section, I look at a model focused on multiple references as the primitive with an additional restriction: the set of references is menu-independent. I provide a characterization of the model and show that in such a model, the set of references can be uniquely identified.
3 Random Attention Model with Unobserved References: Fixed Set of References
In this section, I introduce the reference-dependent random attention model with the restriction that the set of references is menu-independent. Here, there exists a set of alternatives such that whenever an alternative from that set is available in a menu, it is attention privileged.